what is the area of a square plot with side 8 meter ? if it is tiled with square tiles with 50 cm,how much tiles will be needed?
Answer by mananth(16946)

Area of plot = L*W 1m= 100 cm
800*800= 16,0000
Area of tile = 50 * 50 = 2500
Number of tiles = Area of plot/ area of each tile
=160000/2500
= 2025 tiles